# Asymptotic stability of composite waves of two viscous shocks for relaxed compressible Navier-Stokes equations

### v2

@@ 1. Introduction
The constitutive relation (1.1)3italic-(1.1subscriptitalic-)$_{3}$, first introduced by Maxwell in [35], serves to delineate the connection between the stress tensor and velocity gradient for non-simple fluid. The relaxation parameter $\tilde{\tau}=\tilde{\tau}(\rho)$ describe the time lag in response of the stress tensor to velocity gradient. In fact, even for simple fluid, water for example, the time lag does exists but it is very small ranging from 1 ps to 1 ns, see [34, 40]. However, Pelton et al. [38] showed that such a time lag cannot be neglected, even for simple fluids, in the experiments of high-frequency (20GHZ) vibration of nano-scale mechanical devices immersed in water-glycerol mixtures. It was shown that, see also [2], equation (1.1)3italic-(1.1subscriptitalic-)$_{3}$ provides a general formalism with which to characterize the fluid-structure interaction of nano-scale mechanical devies vibrating in simple fluids.
Assuming that $\tilde{\tau}(\rho)=\tau\rho$, where $\tau$ is a positive constant, equation (1.1)3italic-(1.1subscriptitalic-)$_{3}$ reduces to
@@
$\begin{cases}\sigma_{1}(v_{m}-v_{-})=-(u_{m}-u_{-}),\\ \sigma_{1}(u_{m}-u_{-})=p(v_{m})-p(v_{-}),\end{cases}\begin{cases}\sigma_{2}(v_{+}-v_{m})=-(u_{+}-u_{m}),\\ \sigma_{2}(u_{+}-u_{m})=p(v_{+})-p(v_{m}),\end{cases}$ (1.7)
@@
In this paper, we studied the time asymptotic stability for composite waves of two viscous shocks to system (1.3)-(1.4). Note that the dissipation structure of relaxed system (1.3) is much weaker than that of classical system (1.9), thus the BD entropy used in [25, 10] to prove the $a$-contraction property is not available for our system and energy estimates have new challenges. Moreover, compared to the results of [10] where cutoff functions were used to localize the diffusion term, we do not need this kind of operation. We shall follow the methods in [8] to overcome such difficulties. Here are our strategy. Instead of using BD entropy, we use of the special hyperbolic structure of the relaxed system and the relative entropy quantities with weight function and shifts to get the $L^{\infty}_{t}L^{2}_{x}$ estimates of $(v-\widetilde{v},u-\widetilde{u},\Pi-\widetilde{\Pi})$ and weighted $L^{2}_{t}L^{2}_{x}$ estimates of $(v-\widetilde{v},\Pi-\widetilde{\Pi})$ (see Lemma 4.2). We note that, unlike that in [25, 10] where the dissipation estimates of the derivative of $(v-\widetilde{v})$ were obtained for system (1.3), we do not have such estimate. Instead, we shall recover the dissipation estimate of $(v-\widetilde{v})$ from high-order estimates and thus cutoff functions are not needed.
@@
±$\sum\limits_{\pm}\left(\|(v_{0}-v_{\pm},u_{0}-u_{\pm})\|_{L^{2}(\mathbb{R}_{\pm})}\right)+\|((v_{0})_{x},(u_{0})_{x})\|_{H^{1}(\mathbb{R})}+\sqrt{\tau}\|\Pi_{0}\|_{H^{2}(\mathbb{R})}<\varepsilon_{0},$ (1.11)
@@
$(v^{\tau},u^{\tau})\rightharpoonup(v^{0},u^{0})\qquad weakly-*\quad in\quad L^{\infty}\left((0,+\infty);H^{2}\right),$
@@
Notations: $L^{p}(\mathbb{R})$ and $W^{s,p}(\mathbb{R})$ ($1\leq p\leq\infty$) denote the usual Lebesgue and Sobolev spaces over $\mathbb{R}$ with the norm $\|\cdot\|_{L^{p}}$ and $\|\cdot\|_{W^{s,p}}$, respectively. Note that, when $s=0$, $W^{0,p}=L^{p}$. For $p=2$, $W^{s,2}$ are abbreviated to $H^{s}$ as usual. Let $T$ and $B$ be a positive constant and a Banach space, respectively. $C^{k}(0,T;B)(k\geq 0)$ denotes the space of $B$-valued $k$-times continuously differentiable functions on $[0,T]$, and $L^{p}(0,T;B)$ denotes the space of $B$-valued $L^{p}$-functions on $[0,T]$. The corresponding space $B$-valued functions on $[0,\infty)$ are defined in an analogous manner.
@@ 2.1. Traveling wave
Substituting (2.3) and (2.2)1italic-(2.2subscriptitalic-)$_{1}$ into (2.2)3italic-(2.2subscriptitalic-)$_{3}$, we derive that
@@
Let (1.10) hold. For any states $(v_{-},u_{-})$, $(v_{m},u_{m})$, $(v_{+},u_{+})$, and $\sigma_{1}<0,\sigma_{2}>0$ satisfying R-H condition and Lax condition, there exists a positive constant $C$ independent of $\tau$ such that the following is true: the traveling wave solutions $(\widetilde{u}_{1},\widetilde{v}_{1},\widetilde{\Pi}_{1})(\xi_{1})$ connecting $(v_{-},u_{-},0)$ and $(v_{m},u_{m},0)$ and $(\widetilde{u}_{2},\widetilde{v}_{2},\widetilde{\Pi}_{2})(\xi_{2})$ connecting $(v_{m},u_{m},0)$ and $(v_{+},u_{+},0)$ exist uniquely and satisfy
@@
where $P(\widetilde{v}_{1})=\frac{p(v_{-})-p(\widetilde{v}_{1})}{v_{-}-\widetilde{v}_{1}}$. The concavity of $p$ implies
$P^{\prime}(\widetilde{v}_{1})=\frac{-p^{\prime}(\widetilde{v}_{1})(v_{-}-\widetilde{v}_{1})+p(v_{-})-p(\widetilde{v}_{1})}{(v_{-}-\widetilde{v}_{1})^{2}}=\frac{p^{\prime\prime}(v_{\ast})}{2}>0,$
where $v_{\ast}$ is between $v_{-}$ and $v_{m}$. Therefore, we get $\widetilde{v}^{S}_{\xi}<0$.
In addition, using the entropy inequality: $-\sqrt{-p^{\prime}(v_{-})}>-\sqrt{-p^{\prime}(v_{m})}$, we get $v_{-}>v_{m}$. So, we derive that
$v_{m}<\widetilde{v}_{1}<v_{-}.$ (2.6)
Let $f(\widetilde{v}_{1})=\frac{\widetilde{v}_{1}h_{1}(\widetilde{v}_{1})}{\mu\sigma_{1}+\tau\sigma_{1}h^{\prime}_{1}(\widetilde{v}_{1})}$. Using (1.10), one has
$|f(v_{1})-f(v_{2})|$ $=\Big{|}\frac{v_{1}h(v_{1})}{\mu\sigma+\tau\sigma h^{\prime}(v_{1})}-\frac{v_{2}h(v_{2})}{\mu\sigma+\tau\sigma h^{\prime}(v_{2})}\Big{|}$
$\leq\Big{|}\frac{v_{1}h(v_{1})-v_{2}h(v_{2})}{\mu\sigma+\tau\sigma h^{\prime}(v_{1})}\Big{|}+\Big{|}v_{2}h(v_{2})\frac{\sigma\tau\left(h^{\prime}(v_{1})-h^{\prime}(v_{2})\right)}{(\mu\sigma+\tau\sigma h^{\prime}(v_{1}))(\mu\sigma+\tau\sigma h^{\prime}(v_{2}))}\Big{|}$
$\leq C\left(|v_{1}-v_{2}|+|h(v_{1})-h(v_{2})|+|h^{\prime}(v_{1})-h^{\prime}(v_{2})|\right)$
@@
So, $f(\widetilde{v}_{1})$ satisfies the usual Lipschitz condition. This together with (2.6) implies that there exists a unique global solution of (2.5)
@@ 2.2. Construction of shifts
[…]
@@
On the other hand, we known that $21<a<$ 2 and
@@ 3. Local solution and a priori estimates
then the system (1.3) with initial condition (1.4) has a unique solution $(v,u,\Pi)$ on $[0,T_{0}]$ such that
@@
±$\sum\limits_{\pm}\left(\|(\widetilde{v}_{0}-v_{\pm},\widetilde{u}_{0}-u_{\pm},\sqrt{\tau}\widetilde{\Pi}_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)\leq C(\delta_{1}+\delta_{2}).$ (3.4)
The following proposition gives the a priori estimates of the error term: $(v-\widetilde{v},u-\widetilde{u},\Pi-\widetilde{\Pi})$.
@@
Let $(v,u,\Pi)$ be local solutions given by Theorem 3.1 on $[0,T]$ for some $T>0$ and $(\widetilde{v},\widetilde{u},\widetilde{\Pi})$ be defined in (3.1). Then there exist positive constants $\delta_{0},\varepsilon_{1}$ such that if two independent shock waves strength satisfy $\delta_{1},\delta_{2}<\delta_{0}$ and
@@
where $C_{0}$ independent of $T$ and $\tau$ and
$G^{s}(U):=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||v-\widetilde{v}|^{2}\mathrm{d}x$ (3.7)
@@
±$\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{L^{2}(\mathbb{R}_{\pm})}\right)+\|((\underline{v})_{x},(\underline{u})_{x})\|_{H^{1}}<C_{\ast}(\delta_{1}+\delta_{2}).$ (3.9)
@@
±$\leq\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\widetilde{v}_{0}-v_{\pm},\widetilde{u}_{0}-u_{\pm},\sqrt{\tau}\widetilde{\Pi}_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
@@
±$\leq\sum\limits_{\pm}\left(\|(v_{0}-v_{\pm},u_{0}-u_{\pm},\sqrt{\tau}\Pi_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
@@
Therefore, according Theorem 3.1 with (3.12), (3.13) and (3.14), we obtain that the system (1.3) with initial condition (1.4) has a unique solution $(v,u,\Pi)$ on $[0,T_{0}]$ such that
@@
In addition, using the definition of $(\widetilde{v},\widetilde{u},\widetilde{\Pi})$ and Lemma 2.1, we have
@@
±$\leq\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\widetilde{v}-v_{\pm},\widetilde{u}-u_{\pm},\sqrt{\tau}\widetilde{\Pi})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
@@
$X_{1}(t)+\sigma_{1}t\leq\frac{\sigma_{1}}{2}t,\qquad X_{2}(t)+\sigma_{2}t\geq\frac{\sigma_{2}}{2}t,\qquad t>0.$
@@
where $H(v|\widetilde{v})=H(v)-H(\widetilde{v})-H^{\prime}(\widetilde{v})(v-\widetilde{v})$ and potential energy $H(v):=v^{-\gamma+1}/(\gamma-1)$. In addition, for the pressure function $p(v)=v^{-\gamma}$, denote by $p(v|\tilde{v})=p(v)-p(\widetilde{v})-p^{\prime}(\widetilde{v})(v-\widetilde{v})$. Then, we have the following Lemma, see [25, 10].
@@ 4.1. $L^{2}$ energy estimates
Under the hypotheses of Proposition 3.1, there exist $C,C_{2}>0$ (independent of $\tau$ and $T$) such that for all $t\in(0,T]$,
$\|v-\widetilde{v}\|^{2}_{L^{2}}+\|u-\widetilde{u}\|^{2}_{L^{2}}+\tau\|\Pi-\widetilde{\Pi}\|^{2}_{L^{2}}+\int_{0}^{t}\sum\limits_{i=1}\limits^{2}\delta_{i}|\dot{X}_{i}|^{2}\mathrm{d}t+\int_{0}^{t}\left(G^{s}(U)+G(U)\right)\mathrm{d}t$ (4.3)
$\leq C\left(\|v_{0}-\widetilde{v}_{0}(\cdot)\|^{2}_{L^{2}}+\|u_{0}-\widetilde{u}_{0}(\cdot)\|^{2}_{L^{2}}+\tau\|\Pi_{0}-\widetilde{\Pi}_{0}(\cdot)\|^{2}_{L^{2}}\right)+C_{2}\int_{0}^{t}\int_{\mathbb{R}}\left|\partial_{x}(v-\widetilde{v})\right|^{2}\mathrm{d}x\mathrm{d}t+C\delta_{0},$
where
@@
Next, we present the following lemma whose proof parallels the process outlined in Section 4.6 of reference [10]. Despite this methodological resemblance, we have observed that the approach in [10] employs cutoff functions to confine the perturbation, thereby obtaining dissipation estimates, which we have not employed in our proof. To preserve the integrity of our work and to illustrate the distinctions and relationships between our methodology and that of [10], we will provide a detailed exposition of this proof process in the ensuing lemma.
@@
$-\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{2}D$
@@
$D=\int_{\mathbb{R}}\frac{a}{\gamma p(v)}|\partial_{x}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x.$
@@
$w_{1}:=$ $\Big{(}p(v(t,x+X_{1}(t)))-p\left(\widetilde{v}_{1}(x-\sigma_{1}t)+\widetilde{v}_{2}(x-\sigma_{2}t-X_{2}(t)+X_{1}(t))-v_{m}\right)\Big{)}\circ y_{1}^{-1},$ (4.9)
$w_{2}:=$ $\Big{(}p(v(t,x+X_{2}(t)))-p\left(\widetilde{v}_{1}(x-\sigma_{1}t-X_{1}(t)+X_{2}(t))+\widetilde{v}_{2}(x-\sigma_{2}t)-v_{m}\right)\Big{)}\circ y_{2}^{-1}.$
@@
$Y_{11}=\frac{1}{\sigma_{1}^{2}}\int_{\mathbb{R}}a\left(p(\widetilde{v}_{1})^{X_{1}}_{x}-(\widetilde{\Pi}_{1})^{X_{1}}_{x}\right)(p(v)-p(\widetilde{v}))\mathrm{d}x.$
@@
$\frac{1}{\sigma_{1}^{2}}\int_{\mathbb{R}}ap(\widetilde{v}_{1})^{X_{1}}_{x}(p(v)-p(\widetilde{v}))\mathrm{d}x=\frac{\delta_{1}}{\sigma_{1}^{2}}\int_{0}^{1}a(t,x+X_{1}(t))w_{1}\mathrm{d}y_{1}.$
@@
$\Big{|}Y_{11}-\frac{\delta_{1}}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big{|}\leq C\delta_{1}\delta_{0}\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}+C\int_{\mathbb{R}}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x.$
@@
$\Big{|}v-\widetilde{v}-\Big{(}-\frac{1}{\sigma_{m}^{2}}(p(v)-p(\widetilde{v}))\Big{)}\Big{|}\leq C(\delta_{0}+\varepsilon_{1})|p(v)-p(\widetilde{v})|.$
Next, changing the variable $x\rightarrow x+X_{1}(t)$ and applying the new variable for $y_{1},w_{1}$, we derive that
$Y_{12}=-\delta_{1}\int_{0}^{1}a(t,x+X_{1}(t))(v-\widetilde{v})\mathrm{d}y_{1}.$
So, using (4.11) and the definition of $a(t,x)$ in (2.11), the above estimates implies
@@
$\leq C\delta_{1}(\delta_{0}+\varepsilon_{1})\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}.$

### v3

@@ 1. Introduction
[…]
@@
[…]
@@
[…]
@@
$\sum\limits_{\pm}\left(\|(v_{0}-v_{\pm},u_{0}-u_{\pm})\|_{L^{2}(\mathbb{R}_{\pm})}\right)+\|((v_{0})_{x},(u_{0})_{x})\|_{H^{1}(\mathbb{R})}+\sqrt{\tau}\|\Pi_{0}\|_{H^{2}(\mathbb{R})}<\varepsilon_{0},$ (1.11)
@@
$(v^{\tau},u^{\tau})\rightharpoonup(v^{0},u^{0})\qquad weak-*\quad in\quad L^{\infty}\left((0,+\infty);H^{2}\right),$
@@
Notations: $L^{p}(\mathbb{R})$ and $W^{s,p}(\mathbb{R})$ ($1\leq p\leq\infty$) denote the usual Lebesgue and Sobolev spaces over $\mathbb{R}$ with the norm $\|\cdot\|_{L^{p}}$ and $\|\cdot\|_{W^{s,p}}$, respectively. Note that, when $s=0$, $W^{0,p}=L^{p}$. For $p=2$, $W^{s,2}$ are abbreviated to $H^{s}$ as usual. Let $T$ and $B$ be a positive constant and a Banach space, respectively. $C^{k}(0,T;B)(k\geq 0)$ denotes the space of $B$-valued $k$-times continuously differentiable functions on $[0,T]$, and $L^{p}(0,T;B)$ denotes the space of $B$-valued $L^{p}$-functions on $[0,T]$. The corresponding space $B$-valued functions on $0,\infty)$ are defined in an analogous manner.
@@ 2.1. Traveling wave
[…]
@@
Let (1.10) hold. For any states $(v_{-},u_{-}),(v_{m},u_{m}),(v_{+},u_{+})\in\mathbb{R}_{+}\times\mathbb{R}$ and $\sigma_{1}<0,\sigma_{2}>0$ satisfying R-H condition and Lax condition, there exists a positive constant $C$ independent of $\tau$ such that the following is true: the traveling wave solutions $(\widetilde{u}_{1},\widetilde{v}_{1},\widetilde{\Pi}_{1})(\xi_{1})$ connecting $(v_{-},u_{-},0)$ and $(v_{m},u_{m},0)$ and $(\widetilde{u}_{2},\widetilde{v}_{2},\widetilde{\Pi}_{2})(\xi_{2})$ connecting $(v_{m},u_{m},0)$ and $(v_{+},u_{+},0)$ exist uniquely and satisfy
@@
where $P(\widetilde{v}_{1})=\frac{p(v_{-})-p(\widetilde{v}_{1})}{v_{-}-\widetilde{v}_{1}}$.
In addition, using the entropy inequality: $\sqrt{-p^{\prime}(v_{-})}>-\sqrt{-p^{\prime}(v_{m})}$, we get $v_{-}>v_{m}$. So, we derive that
$0<v_{m}<\widetilde{v}_{1}<v_{-}.$ (2.6)
Assuming $\delta_{1}$ is sufficiently small such that $0<\widetilde{v}_{1}-v_{m}<v_{-}-v_{m}<C\delta_{1}<\frac{p^{\prime\prime}(v_{m})}{4}.$ and applying Taylor expansion to the function $P(\widetilde{v}_{1})$ about $v_{m}$, we have
$|P(\widetilde{v}_{1})-P(v_{m})-P^{\prime}(v_{m})(\widetilde{v}_{1}-v_{m})|\leq C|\widetilde{v}_{1}-v_{m}|^{2}\leq C\delta_{1}(\widetilde{v}_{1}-v_{m}).$
Next, the concavity of $p$ implies
$P^{\prime}(v_{m})=\frac{-p^{\prime}(v_{m})(v_{-}-v_{m})+p(v_{-})-p(v_{m})}{(v_{-}-v_{m})^{2}}=\frac{p^{\prime\prime}(v_{\ast})}{2}>0,\quad v_{\ast}\in(v_{m},v_{-}).$
Thus, we get
$P(\widetilde{v}_{1})-P(v_{m})\geq\frac{p^{\prime\prime}(v_{m})}{4}(\widetilde{v}_{1}-v_{m})>0.$
So, we derive that
$h_{1}(\widetilde{v}_{1})>0.$ (2.7)
On the other hand, using (1.8) and (1.10), it holds
$\sigma_{1}(\mu+\tau h^{\prime}_{1}(\widetilde{v}_{1}))<0.$ (2.8)
Therefore, combining (2.6), (2.7), (2.8) and (2.5), we conclude that $(\widetilde{v}_{1})_{\xi_{1}}<0$.
Let $f(\widetilde{v}_{1})=\frac{\widetilde{v}_{1}h_{1}(\widetilde{v}_{1})}{\mu\sigma_{1}+\tau\sigma_{1}h^{\prime}_{1}(\widetilde{v}_{1})}$. For any $v_{1},v_{2}\in(v_{m},v_{-})$, using (1.10), one has
$|f(v_{1})-f(v_{2})|$ $=\Big|\frac{v_{1}h_{1}(v_{1})}{\mu\sigma+\tau\sigma h_{1}^{\prime}(v_{1})}-\frac{v_{2}h_{1}(v_{2})}{\mu\sigma+\tau\sigma h_{1}^{\prime}(v_{2})}\Big|$
$\leq\Big|\frac{v_{1}h_{1}(v_{1})-v_{2}h_{1}(v_{2})}{\mu\sigma+\tau\sigma h_{1}^{\prime}(v_{1})}\Big|+\Big|v_{2}h_{1}(v_{2})\frac{\sigma\tau\left(h_{1}^{\prime}(v_{1})-h_{1}^{\prime}(v_{2})\right)}{(\mu\sigma+\tau\sigma h_{1}^{\prime}(v_{1}))(\mu\sigma+\tau\sigma h_{1}^{\prime}(v_{2}))}\Big|$
$\leq C\left(|v_{1}-v_{2}|+|h_{1}(v_{1})-h_{1}(v_{2})|+|h_{1}^{\prime}(v_{1})-h_{1}^{\prime}(v_{2})|\right)$
@@
So, $f(\widetilde{v}_{1})$ satisfies the usual Lipschitz condition. This together with (2.6) implies that there exists a unique local solution of (2.5)
@@ 2.2. Construction of shifts
$\begin{cases}\begin{aligned} &\dot{X}_{i}(t)=-\frac{M}{\delta_{i}}\left[\int_{\mathbb{R}}\frac{a}{\sigma_{i}}(\widetilde{u}_{i}^{X_{i}})_{x}(p(v)-p(\widetilde{v}))\mathrm{d}x-\int_{\mathbb{R}}a\left(p(\widetilde{v}_{i}^{X_{i}})\right)_{x}(v-\widetilde{v})\mathrm{d}x\right],\\ &X_{i}(t)=0,\end{aligned}\end{cases}$ (2.12)
@@
On the other hand, we known that $1<a<2$ and
@@ 3. Local solution and a priori estimates
then the system (1.3) with initial condition (1.4) has a unique solution $(v,u,\Pi)$ on $0,T_{0}]$ such that
@@
$\sum\limits_{\pm}\left(\|(\widetilde{v}_{0}-v_{\pm},\widetilde{u}_{0}-u_{\pm},\sqrt{\tau}\widetilde{\Pi}_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)\leq C(\delta_{1}+\delta_{2}).$ (3.4)
[…]
@@
[…]
@@
Here $C_{0}$ independent of $T$ and $\tau$ and
$G^{s}(U):=\sum\limits_{i=1}\limits^{2}\int_{\mathbb{R}}|(\widetilde{v}_{i})^{X_{i}}_{x}||\phi_{i}(v-\widetilde{v})|^{2}\mathrm{d}x$ (3.7)
where cutoff functions $\phi_{1},\phi_{2}$ are defined as follows
[…]
$\phi_{2}(x)=1-\phi_{1}(x).$
@@
$\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{L^{2}(\mathbb{R}_{\pm})}\right)+\|((\underline{v})_{x},(\underline{u})_{x})\|_{H^{1}}<C_{\ast}(\delta_{1}+\delta_{2}).$ (3.10)
@@
$\leq\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\widetilde{v}_{0}-v_{\pm},\widetilde{u}_{0}-u_{\pm},\sqrt{\tau}\widetilde{\Pi}_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
@@
$\leq\sum\limits_{\pm}\left(\|(v_{0}-v_{\pm},u_{0}-u_{\pm},\sqrt{\tau}\Pi_{0})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
@@
Therefore, according Theorem 3.1 with (3.13), (3.14) and (3.15), we obtain that the system (1.3) with initial condition (1.4) has a unique solution $(v,u,\Pi)$ on $0,T_{0}]$ such that
@@
[…]
@@
$\leq\sum\limits_{\pm}\left(\|(\underline{v}-v_{\pm},\underline{u}-u_{\pm})\|_{H^{2}(\mathbb{R}_{\pm})}\right)+\sum\limits_{\pm}\left(\|(\widetilde{v}-v_{\pm},\widetilde{u}-u_{\pm},\sqrt{\tau}\widetilde{\Pi})\|_{H^{2}(\mathbb{R}_{\pm})}\right)$
@@
[…]
@@ 4.1. $L^{2}$ energy estimates
Under the hypotheses of Proposition 3.1, there exist constants $C,C_{1}>0$ (independent of $\tau$ and $T$) such that for all $t\in(0,T]$,
$\int_{\mathbb{R}}\eta(U(t,x)|\widetilde{U}(t,x))\mathrm{d}x+\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{4M}\int_{0}^{t}|\dot{X}_{i}|^{2}\mathrm{d}t+C_{1}(1-C(\delta_{0}+\varepsilon_{1}))\int_{0}^{t}G^{s}(U)\mathrm{d}t$ (4.3)
$\quad+(1-\kappa-C(\delta^{\frac{1}{2}}_{1}+\delta^{\frac{1}{2}}_{2}))\int_{0}^{t}G(U)\mathrm{d}t$
$\leq C\int_{\mathbb{R}}\eta(U(0,x)|\widetilde{U}(0,x))\mathrm{d}x+\frac{3\mu}{4v_{m}}\frac{1+\kappa+C\varepsilon_{1}}{1-C\varepsilon_{1}}\int_{0}^{t}\int_{\mathbb{R}}|p^{\prime}(v)||\partial_{x}(v-\widetilde{v})|^{2}\mathrm{d}x\mathrm{d}t+C\delta_{0},$
where $\kappa$ is small constant to be determined later,
@@
Lemma 4.4.
Assume (4.5). Let $\phi_{i}$ be the functions defined in (3.8). Given $v_{+}>0$, there exist positive constants $\delta_{0},C$ such that for any $\delta_{1},\delta_{2}\in(0,\delta_{0})$, the following estimates hold:
$\phi_{2}|(\widetilde{v}_{1})_{x}^{X_{1}}|\leq C\delta_{1}^{2}e^{-C\delta_{1}t},\quad\phi_{1}|(\widetilde{v}_{2})_{x}^{X_{2}}|\leq C\delta_{2}^{2}e^{-C\delta_{2}t}\quad t>0,\quad x\in\mathbb{R},$
$\int_{\mathbb{R}}\phi_{2}|(\widetilde{v}_{1})_{x}^{X_{1}}|\mathrm{d}x\leq C\delta_{1}e^{-C\delta_{1}t},\quad\int_{\mathbb{R}}\phi_{1}|(\widetilde{v}_{2})_{x}^{X_{2}}|\mathrm{d}x\leq C\delta_{2}e^{-C\delta_{2}t}\quad t>0.$
By Lemma 2.1 and the definition of $\phi_{i}$, this lemma follows immediately. we omit the proof and refer to Appendix B of [8] for details.
@@
Next, we establish the following lemma. Our proof is inspired by the core idea of using cutoff functions, as introduced in [8], to manage perturbations. Nonetheless, we have developed a refined set of estimates and a modified technical approach to suit our specific framework. The following self-contained proof is provided to clearly delineate our methodology and its departures from the source.
@@
$-\sum\limits_{i=1}\limits^{2}\frac{\delta_{i}}{2M}|\dot{X}_{i}|^{2}+B_{1}+B_{2}-G_{2}-\frac{3}{4}D$
@@
$\quad+C\left(\sum\limits_{i=1}\limits^{2}\delta_{i}^{2}e^{-C\delta_{i}t}+\frac{1}{t^{2}}\right)\int_{\mathbb{R}}\eta(U|\widetilde{U})\mathrm{d}x,$
@@
$D=\int_{\mathbb{R}}\frac{a\mu}{\gamma p(v)}|\partial_{x}(p(v)-p(\widetilde{v}))|^{2}\mathrm{d}x.$
@@
$w_{1}$= $\phi_{1}(x+X_{1}(t))\Big(p(v(t,x+X_{1}(t)))$ (4.9)
$-p\left(\widetilde{v}_{1}(x-\sigma_{1}t)+\widetilde{v}_{2}(x-\sigma_{2}t-X_{2}(t)+X_{1}(t))-v_{m}\right)\Big)\circ y_{1}^{-1},$
$w_{2}$= $\phi_{2}(x+X_{2}(t))\Big(p(v(t,x+X_{2}(t)))$
$-p\left(\widetilde{v}_{1}(x-\sigma_{1}t-X_{1}(t)+X_{2}(t))+\widetilde{v}_{2}(x-\sigma_{2}t)-v_{m}\right)\Big)\circ y_{2}^{-1}.$
@@
$Y_{11}=\frac{1}{\sigma_{1}^{2}}\int_{\mathbb{R}}a\phi_{1}\left(p(\widetilde{v}_{1})^{X_{1}}_{x}-(\widetilde{\Pi}_{1})^{X_{1}}_{x}\right)(p(v)-p(\widetilde{v}))\mathrm{d}x+\int_{\mathbb{R}}a\phi_{2}(\widetilde{v}_{1})^{X_{1}}_{x}(p(v)-p(\widetilde{v}))\mathrm{d}x.$
@@
$\frac{1}{\sigma_{1}^{2}}\int_{\mathbb{R}}a\phi_{1}p(\widetilde{v}_{1})^{X_{1}}_{x}(p(v)-p(\widetilde{v}))\mathrm{d}x=\frac{\delta_{1}}{\sigma_{1}^{2}}\int_{0}^{1}a(t,x+X_{1}(t))w_{1}\mathrm{d}y_{1}.$
@@
$\Big|Y_{11}-\frac{\delta_{1}}{\sigma_{m}^{2}}\int_{0}^{1}w_{1}\mathrm{d}y_{1}\Big|\leq C\delta_{1}\delta_{0}\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}+C\int_{\mathbb{R}}\Big(\phi_{2}|(\widetilde{v}_{1})^{X_{1}}_{x}|+\phi_{1}|(\widetilde{\Pi}_{1})^{X_{1}}_{x}|\Big)|p(v)-p(\widetilde{v})|\mathrm{d}x.$
@@
$p(v~)|.\Big|v-\widetilde{v}-\Big(-\frac{1}{\sigma_{m}^{2}}(p(v)-p(\widetilde{v}))\Big)\Big|\leq C(\delta_{0}+\varepsilon_{1})$|p(v)-p($\widetilde${v})|.
Next, changing the variable $x\rightarrow x+X_{1}(t)$, applying the new variable for $y_{1},w_{1}$ and using (4.11) and the definition of $a(t,x)$ in (2.13), it holds
@@
$\quad+C\int_{\mathbb{R}}\phi_{2}|(\widetilde{v}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x$
$\leq C\delta_{1}(\delta_{0}+\varepsilon_{1})\int_{0}^{1}|w_{1}|\mathrm{d}y_{1}+C\int_{\mathbb{R}}\phi_{2}|(\widetilde{v}_{1})^{X_{1}}_{x}||p(v)-p(\widetilde{v})|\mathrm{d}x.$